OpenStudy (anonymous):

Were finding the slope intercept form of the equation.. i need help on, -13/6=-1/6y + 1x if anyone understand how to do this please help meee;p thakns! :)

OpenStudy (anonymous):

1/6y=x+13/6 Cross multiplying, y=6x+13. which is the slope intercept form.slope is 6 and y intercept is 13

OpenStudy (anonymous):

thank you!

OpenStudy (ash2326):

@powers98 Please don't post Maths questions in Physics. This section is only for Physics. Thanks:)

OpenStudy (anonymous):

Im sorry! i didnt realize this was the wrong one!

OpenStudy (ash2326):

It's ok but remember to post in the right group next time